# Experimental Demonstration of a Quantum-Optimal Coronagraph Using Spatial Mode Sorters

### v1

@@ Abstract
An ideal direct imaging coronagraph, which selectively rejects the fundamental mode of a telescope, has been shown to achieve the quantum information limits for exoplanet detection and localization. In this study, we experimentally implement this quantum-optimal coronagraph using spatial mode (de)multiplexing. Our benchtop system includes a forward and inverse pass through a free-space programmable spatial mode sorter, designed to isolate photons in a point spread function (PSF)-adapted basis. During the forward pass, the fundamental mode is rejected, effectively eliminating light from an on-axis point-like star. On the inverse pass, the remaining modes are coherently recombined, enabling direct imaging of a faint companion. We develop a probabilistic measurement model that accounts for combined effects of fundamental shot noise and experimental noise specific to our benchtop setup, such as modal cross-talk, dark noise, and ambient background illumination. We leverage this measurement model to formulate a maximum-likelihood estimator of the exoplanet position given an image captured with the coronagraph. Using this approach, we successfully localize an artificial exoplanet at sub-diffraction distances $(<\sigma)$ from its host star under a 1000:1 star-planet contrast ratio. Our system accurately localizes the exoplanet up to an absolute error $<0.03\sigma$ over the separation range [0, $[0,\,0.6]\sigma$. Finally, we numerically evaluate the precision of our experimental coronagraph against state-of-the-art coronagraphs subject to comparable noise models.
@@ I Introduction
By comparison, direct imaging techniques aspire to spatially observe/resolve orbiting exoplanets, providing more comprehensive planetary data Seager and Deming (2010); Biller and Bonnefoy (2018). However, direct imaging faces two compounding phenomenological challenges. First, exoplanets are extremely faint compared to their host stars, with relative brightness factors ranging from $10^{-5}$ for Hot Jupiters to $10^{-11}$ for Exo-Earths in the habitable zone. Second, the distance between an exoplanet and its host star often falls below the optical resolution capabilities of current space-based telescopes, residing in the so-called ’sub-diffraction regime’ Biller (2013). When imaged with a conventional telescope, light from the exoplanet overlaps with prominent diffraction features of the host star. This overlap, combined with the overwhelming shot noise generated by the bright star, effectively renders the exoplanet undetectable. Developments in coronagraph techniques have enabled complete nulling of an on-axis point-like star so that, in the absence of ambient background illumination, only light from the exoplanet reaches the detector Foo et al. (2005); Mari et al. (2012); Guyon et al. (2010); Aime et al. (2002); Soummer et al. (2003); Soummer (2004). In this way, state-of-the-art coronagraphs suppress photon shot noise intrinsic to measurements of classical states of light, thereby enhancing the signal-to-noise ratio of exoplanet signatures Zurlo (2024); Guyon et al. (2006).
@@
Figure 1: Comparisons between the Chernoff exponent (a) and the Fisher information per photon (b) for various coronagraph designs relative to the quantum limits (black) of exoplanet detection and localization at high contrasts. The Perfect Coronagraph (dashed yellow) is the focus of this work and is shown to theoretically saturate both quantum limits in Deshler et al. (2024). State-of-the-art systems such as the phase-induced amplitude apodization complex mask coronagraph (PIAACMC) Guyon et al. (2010) and the vortex coronagraph Foo et al. (2005) are notably sub-optimal in the sub-diffraction regime.
@@
Inspired by new insights in passive superresolution imaging Tsang et al. (2016); Tsang (2019); Lee et al. (2023); Prasad (2020a), we recently reported the quantum information limits for exoplanet detection and localization Deshler et al. (2024). Our findings revealed that these limits are achieved by a direct-imaging coronagraph that exclusively rejects the fundamental mode of the telescope. In contrast, current state-of-the-art coronagraphs discard information-bearing photons in higher-order modes Belikov et al. (2021), resulting in sub-optimal performance over the sub-diffraction regime, as illustrated in Figure 1. Quantum-optimal coronagraphs, however, preserve information at sub-diffraction star-planet separations, where an abundance of exoplanets are expected to reside, given current statistical models Lagrange et al. (2023); Fernandes et al. (2019); Chen and Kipping (2016); Ning et al. (2018).
@@
In this work, we propose a quantum-optimal direct imaging coronagraph using a spatial mode sorter implemented with a multi-plane light converter (MPLC) Carpenter et al. (2020); Fontaine et al. (2019); Labroille et al. (2014). To the best of our knowledge, this is the first experimental verification of a coronagraph design that theoretically saturates the quantum limits for exoplanet discovery tasks. Applying a maximum likelihood estimator to images collected with our bench-top setup, we localize an artificial exoplanet at sub-diffraction separations from an artificial host star with a contrast ratio of 1000:1. Denoting the Rayleigh diffraction limit of our imaging system as $\sigma$, the absolute error of the empirical mean MLE remains below $0.03\sigma$ over the separation range [0, $[0,\,0.6]\sigma$. The empirical precision (standard deviation) of the MLE varies over the sub-diffraction regime between $\sim 0.1\sigma$ and $\sim 0.01\sigma$ for an exoplanet in the range $[0,0.1]\sigma$ and $[0.1,0.6]\sigma$, respectively. We invoke a probabilistic measurement model to characterize the impact of real-world noise sources and experimental constraints expected in an operating environment which hinder the attainment of quantum-limited performance set by shot noise.
@@ II Methods
Figure 2: (a) Conceptual design of a quantum-optimal direct-imaging coronagraph based on spatial mode sorting. The forward mode sorter demultiplexes the field incident on the image plane into a PSF-adapted basis. Light in the fundamental mode is rejected by a mask at the sorting plane while light in the remaining modes propagates freely to an inverse mode sorter. The inverse mode sorter coherently recombines the light to form an image at the detector. (b) Schematic of the experimental implementation of a quantum-optimal coronagraph using an MPLC mode sorter and non-reciprocal polarization elements. (c) Intensity profiles of the truncated Fourier-Zernike mode basis used in our experiment.
@@ II.1 Experimental Design
Figure 2(a) illustrates the working principles of a quantum-optimal coronagraph design based on two cascaded mode sorters. The first mode sorter decomposes the incident optical field into a PSF-adapted transverse spatial mode basis Řehaček et al. (2017) in order to isolate and eliminate photons in the fundamental mode. The second sorter inverts the mode decomposition, coherently recombining light in the residual modes to form an image of the exoplanet on a detector array. This scheme can be viewed as spatial mode filtering.
@@
Our experimental setup shown in Figure 2b emulates these working principles by double-passing the optical field through a single mode sorter implemented on a 3-plane MPLC Ozer et al. (2022). On the forward pass, the MPLC spatially demultiplexes the optical field in the Fourier-Zernike modes (Appendix A), which constitute a PSF-adapted basis for circular apertures, and focuses light in each mode to a distinct Gaussian spot on the sorting plane. The spot corresponding to the fundamental mode is directed to the opening of a pinhole mirror and absorbed at a beam dump. The remaining modes reflect off the pinhole mirror and are sent backwards through the mode sorter. The unitary nature of spatial mode sorting inverts the mode transformation during the backward pass. Non-reciprocal polarization elements split the optical path for the forward (pre-nulling) and backward (post-nulling) pass, sending the filtered field to a detector. Through this process, the field at the detector plane is identical to the field at the focal plane minus optical contribution from the fundamental mode. The $4f$ imaging system used in our setup is characterized by a circular aperture diameter $D=400\mu\text{m}$ and a focal length $f=200\text{mm}$ operating at wavelength $\lambda=532\text{nm}$, yielding a Rayleigh resolution of $\sigma=1.22\lambda f/D=324\mu\text{m}$ on the object plane.
@@
We prioritize demonstrating exoplanet localization at sub-Rayleigh star-planet separations; the regime where quantum-optimal coronagraphs offer the greatest theoretical advantage over existing high-performance coronagraph designs. To sample this separation regime, we align the coronagraph to a bright on-axis point-source (artificial star) and vertically step the position of a second dim point-source (artificial exoplanet) $\vec{r}_{e}=(0,y_{e})$ over the discrete domain $y_{e}\in\mathcal{Y}=[-a:\Delta:+a]$ with endpoints $a=.85\sigma$ and sampling step size $\Delta=.0215\sigma$.
@@
Light from a sub-diffraction exoplanet couples predominantly to lower-order Fourier-Zernike modes. We therefore configured the MPLC to sort a truncated basis $\{\bm{\psi}_{0}(\vec{r}),\bm{\psi}_{1}(\vec{r}),\bm{\psi}_{2}(\vec{r}),\bm{\psi}_{3}(\vec{r})\}$ where $\bm{\psi}_{0}(\vec{r})$ is the fundamental mode of the imaging system (Figure 2c). Collectively, these modes contain a majority of the energy in the field generated by a sub-Rayleigh companion as shown in Figure 3. We define the nominal region of support for this truncated basis to be $|y_{e}|\leq.6\sigma$. Expanding the basis to more than four modes was found to significantly degrade the cross-talk of the mode sorter due to the limited number of phase masks available on our programmable MPLC. In principle, introducing more masks would allow one to sort more modes and temper modal cross-talk, enabling access to greater star-planet separations and contrasts.
@@
Figure 3: (a) Theoretical mode detection probabilities as a function of off-axis source location for the truncated set of Fourier-Zernike modes. The nominal ’region of support’ is given by off-axis positions in the domain $[0,0.6]\sigma$ within which $>95\%$ of the light energy couples to the truncated mode basis. (b) Distribution of the classical Fisher information in each mode as a function of off-axis position. The classical Fisher information of the truncated mode basis remains within $10\%$ of the quantum Fisher information limit over the domain $\sim[0,0.4]\sigma$. Outside this domain, the information concentrates in higher-order modes that are ignored by our mode-sorter.
@@ II.2 Measurement Model
Figure 4: (a) Cross-talk calibration data (dotted) of relative mode intensity in each channel as a function of off-axis source position. The relative intensities were measured at the sorting plane of the MPLC. Solid curves correspond to theoretically predicted mode intensities under a least-squares fit of the unitary cross-talk matrix to calibration data. (b) Magnitude of the entries in the cross-talk matrix. The relative power leakage from the fundamental mode into all other modes (square of column 1 entries) sums to $\sim 10^{-3}$, setting the limit of accessible star-planet contrasts for our experimental setup. (c) Measurements at the sorting plane with region segmentation for measuring energy in each mode.
We invoke a theoretically and empirically-driven probability distribution for the direct imaging measurement. Let $\mathbf{X}(\vec{r}_{0})\in\mathbb{R}^{M}$ be a random vector containing the number of photons measured at each pixel of the detector over fixed exposure time $T$ when imaging a single point-source located at position $\vec{r}_{0}$. The number of photons measured in each pixel are independent and modeled with Poisson statistics,
@@
where $\lambda_{0}\in\mathbb{R}$ is the photon flux entering the pupil from the point source, $\mathbf{q}(\vec{r}_{0})\in\mathbb{R}^{M}$ is the post-nulling photon arrival probability at each detector pixel after propagating the pupil field through the coronagraph, and $\mathbf{p}_{B}\in\mathbb{R}^{M}$ is an experimentally-observed background distribution with flux rate $\lambda_{B}\in\mathbb{R}$ (Appendix B). For simplicity, the flux rates $\lambda_{0},\lambda_{B}$ are given in photons per integration period $T$. The action of the coronagraph on the incoming optical field appears in the post-nulling photon arrival probability,
$\mathbf{q}(\vec{r}_{0})=|\Psi\Omega C\Omega^{\dagger}\Psi^{\dagger}\psi_{0}(\vec{r}_{0})|^{2},$ (2)
where the input field $\psi_{0}(\vec{r}_{0})$ is a shifted version of the fundamental mode induced by illumination by a point source located at $\vec{r}_{0}$. The $|\cdot|^{2}$ operation is applied element-wise to convert the field at each pixel to intensity. We have also introduced several system-dependent matrices: $\Psi\in\mathbb{C}^{M\times K}$ is a truncated change-of-basis matrix that transforms a field from its modal representation to its spatial representation, $C\in\mathbb{R}^{K\times K}$ is a diagonal nulling matrix which acts to reject the fundamental mode, and $\Omega\in\mathbb{C}^{K\times K}$ is the unitary cross-talk matrix of the mode sorter whose entries are determined from calibration measurements (Appendix C).
A synthetic measurement of a star-planet system $\mathbf{Y}=\mathbf{X}_{s}+\mathbf{X}_{e}$ is constructed by adding multiple measurement realizations of the artificial star and exoplanet illuminated independently such that $\mathbf{X}_{s}=\sum_{i=1}^{N_{s}}\mathbf{X}^{(i)}(\vec{0})$ and $\mathbf{X}_{e}=\mathbf{X}(\vec{r}_{e})$ where $\vec{r}_{e}$ is the position of the exoplanet. The choice of $N_{s}$ sets the star-planet contrast. We employ this synthetic measurement scheme to circumvent interference effects that would arise if both star and exoplanet were illuminated simultaneously by the laser source in our setup. The synthetic measurement thus constitutes an approximation of measuring two incoherent point sources with unequal brightness. The complete measurement model is given by,
@@
where $\Lambda_{0}=N\lambda_{0}$, $\Lambda_{B}=N\lambda_{B}$, and $N=N_{s}+1$ being the total number of measurement realizations. We define the photon distribution from the star-planet system as,

### v3

@@ Abstract
Deep sub-diffraction exoplanet discovery currently lies beyond the reach of state-of-the-art direct imaging coronagraphs, which typically have an inner working angle larger than the diffraction scale. We present an experimental demonstration of a direct imaging coronagraph design capable of achieving the quantum limits of exoplanet detection and localization below the Rayleigh diffraction limit. Our benchtop implementation performs a forward and inverse pass through a free-space programmable spatial mode sorter configured to isolate photons in a point spread function (PSF)-adapted mode basis. During the forward pass, the fundamental mode is rejected, effectively eliminating light from an on-axis point-like star. On the inverse pass, the remaining modes are coherently recombined to form an image of a faint companion. Our experimental system is shown localizing an artificial exoplanet at sub-diffraction distances from its host star under a 1000:1 star-planet contrast.
@@ I Introduction
By comparison, direct imaging techniques aspire to spatially observe/resolve orbiting exoplanets, providing more comprehensive planetary data Seager and Deming (2010); Biller and Bonnefoy (2018). However, direct imaging faces two compounding obstacles: photon shot noise and diffraction. Exoplanets are extremely faint compared to their host stars, with relative brightness factors ranging from $10^{-5}$ for Hot Jupiters to $10^{-11}$ for Exo-Earths in the habitable zone. Moreover, the distance between an exoplanet and its host star often falls below the optical resolution capabilities of current space-based telescopes, residing in the so-called ‘sub-diffraction regime’ Biller (2013). Under these circumstances, light from the exoplanet overlaps with prominent diffraction features of the host star. This overlap, combined with the overwhelming shot noise generated by the bright star, effectively renders the exoplanet undetectable with conventional telescopes. Developments in coronagraphy techniques have succeeded in nulling an on-axis point-like star so that only light from the exoplanet reaches the detector Foo et al. (2005); Mari et al. (2012); Guyon et al. (2010); Aime et al. (2002); Soummer et al. (2003); Soummer (2004). In this way, state-of-the-art coronagraphs suppress photon shot noise intrinsic to measuring classical states of light, thereby enhancing the signal-to-noise ratio (SNR) of exoplanet signatures Zurlo (2024); Guyon et al. (2006).
@@
Figure 1: Comparisons between the Chernoff exponent (a) and the Fisher information per photon (b) as a function of the normalized star-planet separation $r_{\Delta}/\sigma$ for various coronagraph designs relative to the quantum limits (black) of exoplanet detection and localization at high contrasts. The Perfect Coronagraph (dashed yellow) is the focus of this work and is known to theoretically saturate both quantum limits in Deshler et al. (2024). State-of-the-art systems such as the phase-induced amplitude apodization complex mask coronagraph (PIAACMC) Guyon et al. (2010) and the vortex coronagraph Foo et al. (2005) are notably sub-optimal in the sub-diffraction regime. Spatial mode demultiplexing (SPADE) Tsang et al. (2016) is an alternative quantum-optimal measurement shown for reference, though it does not qualify as a direct imaging coronagraph.
@@
Inspired by new insights in passive super-resolution imaging Tsang et al. (2016); Tsang (2019a); Lee et al. (2023); Prasad (2020a), we recently reported the quantum information limits for exoplanet detection and localization Deshler et al. (2024). Our findings revealed that these limits are achieved by a direct-imaging coronagraph that exclusively rejects the fundamental mode of the telescope. In contrast, current state-of-the-art coronagraphs discard information-bearing photons in higher-order modes Belikov et al. (2021), resulting in sub-optimal performance over the sub-diffraction regime, as illustrated in Figure 1. Quantum-optimal coronagraphs, however, preserve information at sub-diffraction star-planet separations, where an abundance of exoplanets are expected to reside given current statistical models Lagrange et al. (2023); Fernandes et al. (2019); Chen and Kipping (2016); Ning et al. (2018).
Emerging spatial mode demultiplexing (SPADE) coronagraph platforms such as photonic lanterns Xin et al. (2022, 2023) offer an alternative sensing paradigm that involves measuring optical power in different spatial modes. Although this approach provides a promising direction for exoplanet spectroscopy, it departs from the imaging paradigm, which we argue remains desirable in coronagraphy. In brief, images immediately provide the context and composition of a scene, thereby relaxing the amount of prior information needed to interpret and process measurements. Consider, for example, a prospective host star surrounded by multiple luminous sources (e.g. systems of exoplanets or exozodiacal dust). The distribution of these sources around the star is impossible to uniquely determine from modal power measurements alone, without additionally measuring the relative phase between modes or sorting multiple mode bases Kim et al. (2024); Tsang (2019b, 2021). In imaging, such information is more readily apparent.
@@
In this work, we propose a direct imaging coronagraph using a spatial mode sorter implemented with a multi-plane light converter (MPLC) Carpenter et al. (2020); Fontaine et al. (2019); Labroille et al. (2014). To the best of our knowledge, this is the first experimental verification of a coronagraph design that theoretically saturates the quantum limits for exoplanet discovery tasks. Applying a maximum likelihood estimator (MLE) to the images collected with our benchtop setup, we localize an artificial exoplanet at sub-diffraction separations from an artificial host star with a contrast ratio of 1000:1. Denoting the Rayleigh diffraction limit of our imaging system as $\sigma$, our experiment achieves an absolute localization error below $0.03\sigma$ over the separation range [0, $[0,\,0.6]\sigma$ with the MLE.
@@ II Methods
Figure 2: (a) Conceptual design of a quantum-optimal direct-imaging coronagraph based on spatial mode sorting. The forward mode sorter demultiplexes the field incident on the image plane into a PSF-adapted basis. Light in the fundamental mode is rejected by a mask at the sorting plane while light in the remaining modes propagates freely to an inverse mode sorter. The inverse mode sorter coherently recombines (multiplexes) the remaining modes to form an image at the detector. (b) A simplified schematic of the experimental implementation of a quantum-optimal coronagraph using an MPLC mode sorter and non-reciprocal polarization elements. The MPLC is comprised of a sequence of diffractive phase elements designed to (de)multiplex Fourier-Zernike modes $\{\bm{\psi}_{0},\bm{\psi}_{1},\bm{\psi}_{2},\bm{\psi}_{3}\}$ where $\bm{\psi}_{0}$ is the PSF of the imaging system. A maximum likelihood estimator is applied on the captured image to localize the position of a sub-diffraction exoplanet.
@@ II.1 Experimental Design
Figure 2(a) illustrates the working principles of a quantum-optimal coronagraph design based on two cascaded mode sorters. The first mode sorter decomposes the incident optical field into a PSF-adapted transverse spatial mode basis Řehaček et al. (2017) in order to isolate and eliminate photons residing in the fundamental mode. The second sorter inverts the mode decomposition, coherently recombining light in the residual modes to form an image of the exoplanet on a detector array. This scheme can be viewed as spatial mode filtering.
@@
Our experimental setup shown in Figure 2(b) emulates the working principles of Figure 2(a) by double-passing the optical field through a single mode sorter implemented on a 3-plane MPLC detailed in Ozer et al. (2022, 2024). During the forward pass, the MPLC spatially demultiplexes the optical field in the Fourier-Zernike modes $\{\bm{\psi}_{nm}(\vec{r})\}$ (defined in Appendix A), which constitute a PSF-adapted basis for circular apertures, and focuses light in each mode to spatially resolved Gaussian spots on the sorting plane. The spot corresponding to the fundamental mode is directed to the opening of a pinhole mirror and absorbed at a beam dump. The remaining modes reflect off the pinhole mirror and are sent backwards through the mode sorter. The unitary nature of spatial mode sorting reverses the mode transformation during the backward pass. Non-reciprocal polarization elements split the optical path for the forward (pre-nulling) and backward (post-nulling) pass, sending the filtered field to a detector. Through this process, the field at the detector plane is identical to the field at the focal plane minus optical contribution from the fundamental mode. The $4f$ imaging system used in our setup is characterized by a circular aperture diameter $D=400\mu\text{m}$ and a focal length $f=200\text{mm}$ operating at wavelength $\lambda=532\text{nm}$, yielding a Rayleigh resolution of $\sigma=1.22\lambda f/D=324\mu\text{m}$ on the object plane.
@@
We focus on demonstrating exoplanet localization at sub-Rayleigh star-planet separations; the regime where quantum-optimal coronagraphs offer the greatest theoretical advantage over existing high-performance coronagraph designs. To sample this separation regime, we align the coronagraph to a bright on-axis point-source (artificial star) and vertically step the position of a second dim point-source (artificial exoplanet) $\vec{r}_{e}=(0,y_{e})$ over the discrete domain $y_{e}\in\mathcal{Y}=[-h:\Delta:+h]$ with endpoints $h=0.85\sigma$ and sampling step size $\Delta=0.0215\sigma$.
@@
Light from a sub-diffraction exoplanet couples predominantly to a small set of lower-order Fourier-Zernike modes. We therefore configured the MPLC to sort a truncated basis $\{\bm{\psi}_{0},\bm{\psi}_{1},\bm{\psi}_{2},\bm{\psi}_{3}\}=\{\bm{\psi}_{00},\bm{\psi}_{1-1},\bm{\psi}_{20},\bm{\psi}_{22}\}$ depicted in Figure 2(b) where $\bm{\psi}_{0}(\vec{r})$ is the fundamental mode of the imaging system. Collectively, these modes contain a majority of the energy in the field generated by a sub-Rayleigh companion traversing the $y$-axis as shown in Figure 3(a). We define the nominal region of support for this truncated basis to be $|y_{e}|\leq 0.6\sigma$. For any exoplanet located in this region, >95% of the optical energy couples to the truncated mode basis.
The mode set used in our experimental setup was chosen specifically for 1D localization of a sub-diffraction exoplanet displaced along the y-axis. However, it is straightforward to configure the setup for 2D localization by augmenting/modifying the truncated basis to include modes which are sensitive to displacement along the x-axis. For example, we could have chosen to sort the modes $\{\bm{\psi}_{00},\bm{\psi}_{1-1},\bm{\psi}_{11},\bm{\psi}_{20}\}$ which affords sub-diffraction sensitivity to horizontal displacement through the $\bm{\psi}_{11}$ mode. Expanding the basis to more than four modes was found to significantly degrade the cross-talk of the mode sorter due to the limited number of phase masks available on our programmable MPLC. In principle, introducing more masks would allow one to sort more modes and temper modal cross-talk, enabling access to larger star-planet separations and higher contrasts.
@@
Figure 3: (a) Theoretical photon detection probabilities (coupling strengths) in each mode as a function of off-axis source location for the truncated set of Fourier-Zernike modes in the absence of cross-talk. The nominal ’region of support’ is given by off-axis positions in the domain $[0,0.6]\sigma$ within which $>95\%$ of the light energy couples to the truncated mode basis. (b) Distribution of the classical Fisher information in each mode as a function of off-axis position. The classical Fisher information of the truncated mode basis remains within $10\%$ of the quantum Fisher information limit over the domain $\sim[0,0.4]\sigma$. Outside this domain, the information concentrates in higher-order modes that are ignored by our mode-sorter.
@@ II.2 Measurement Model
Figure 4: (a) Cross-talk calibration data (dotted) depicting the relative optical intensity measured in each target region of the sorting plane as a function of off-axis source position. Solid curves correspond to theoretically predicted mode intensities under a constrained least-squares fit of the unitary cross-talk matrix to the calibration data. (b) A visual representation of the least-squares fitting algorithm used to estimate the experimental cross-talk matrix under the constraint that $\Omega$ be a unitary matrix (further details in Appendix D). (c) Example calibration images of the optical intensity distribution measured at the sorting plane for various off-axis source locations. The total intensity of each mode was taken to be the integrated intensity within their designated circular region on the detector.
Figure 5: (a) Amplitude and (b) phase of the experimental cross-talk matrix recovered from the best-fit calibration method (depicted in Figure 4) under the approximation that $\Omega$ be unitary. This approximation ignores scattering into modes beyond the span of the truncated mode set (see Appendix E for details).The first row of the amplitude matrix corresponds to leakage from the fundamental mode into higher order modes (imperfect star nulling), while the first column of the amplitude matrix corresponds to leakage from the higher order modes into the fundamental mode (lost exoplanet light). The phase components of the experimental cross-talk matrix are artifacts of the Fourier-Zernike mode definitions and the unitary approximation of $\Omega$.
Figure 6: End-to-end throughput of the coronagraph as measured at the image plane (black scatter dots). The throughput predicted by the experimental model (solid black curve) makes use of the MPLC cross-talk matrix of Figure 5. For reference, we also plot the ideal throughput (solid red curve) achieved by a perfect coronagraph that exclusively rejects the fundamental mode. The inset depicts the throughput on log-scale near the optical axis. When the source is on-axis, the measured throughput approximates the leakage from the fundamental mode predicted by the cross-talk matrix $\sim 10^{-3}$.
Figure 7: A comparison between theoretical, experimental, and simulated direct imaging measurements of the artificial exoplanet for various sub-diffraction star-planet separations along the y-axis under a 1000:1 star-planet contrast. In each image, the ground truth location of the exoplanet coincides with the white cross-hairs while the star is located at the origin. All images are max-normalized to illustrate dominant spatial features. THEORETICAL: These images correspond to the perfect coronagraph operating in the absence of background or modal cross-talk. EXPERIMENTAL: These are images of the artificial exoplanet collected with our bench top coronagraph after subtraction of the background. SIMULATED: These images were generated by numerically propagating the optical field through a digital twin of the MPLC device used in our experimental setup. Noticeable distortions and asymmetries in the experimental and simulated images for equidistant off-axis ±$\pm y_{e}$ exoplanet locations arise from modal cross-talk matrices depicted in Figure 5 and Figure 14 respectively.
We construct a statistical forward model for the images captured with our experimental coronagraph. To circumvent erroneous interference effects that would arise if both star and exoplanet were simultaneously illuminated by the coherent laser source in our setup, we synthesize a measurement of the star-planet scene by adding measurements of each body illuminated independently, $\mathbf{Y}=\mathbf{X}_{s}+\mathbf{X}_{e}$. Here $\mathbf{X}_{s}$ and $\mathbf{X}_{e}$ are random vectors denoting image captured with our coronagraph when illuminating the artificial star and exoplanet respectively. Importantly, the statistics of the synthetic measurement $\mathbf{Y}$ are identical to that expected for an incoherent star-planet pair.
The $\mathbf{X}_{s}$ and $\mathbf{X}_{e}$ are themselves constructed from independent realizations of a single-shot measurement. Specifically, $\mathbf{X}_{s}=\sum_{i=1}^{N_{s}}\mathbf{X}^{(i)}(\vec{0})$ and $\mathbf{X}_{e}=\mathbf{X}(\vec{r}_{e})$ where $\vec{r}_{e}$ is the position of the exoplanet and the choice of $N_{s}$ sets the star-planet contrast. Meanwhile, the single-shot measurement $\mathbf{X}(\vec{r}_{0})\in\mathbb{N}^{M}$ is a random vector denoting the number of photons measured at each pixel of a $M$-pixel detector array over fixed exposure time $T$ when imaging a single point-source located at position $\vec{r}_{0}$. The number of photons measured in each pixel are statistically independent Poisson random variables such that,
@@
where $\lambda_{0}\in\mathbb{R}$ is the photon flux entering the pupil from the point source, $\mathbf{q}(\vec{r}_{0})\in\mathbb{R}^{M}$ is the post-nulling photon arrival probability at each pixel after propagating the pupil field through the coronagraph, and $\mathbf{p}_{B}\in\mathbb{R}^{M}$ is an experimentally-observed optical background distribution with flux rate $\lambda_{B}\in\mathbb{R}$ (Appendix B). For simplicity, the flux rates $\lambda_{0}$ and $\lambda_{B}$ are expressed in photons per integration period $T$. The effect of the coronagraph on the incoming optical field manifests in the post-nulling photon arrival probability,
$\mathbf{q}(\vec{r}_{0})=|\Psi\Omega C\Omega^{\dagger}\Psi^{\dagger}\mathbf{u}(\vec{r}_{0})|^{2},$ (2)
where the input field $\mathbf{u}(\vec{r}_{0})=\text{vec}[\bm{\psi}_{0}(\vec{r}-\vec{r}_{0})]\in\mathbb{C}^{M}$ is a square-normalized discrete version of the shifted PSF induced by illumination from a point source located at $\vec{r}_{0}$ (Appendix C). The $|\cdot|^{2}$ operation is applied element-wise to convert the field at each pixel to intensity. We have introduced several system-dependent matrices that map between the image space of dimension $M$ and a truncated mode space of dimension $K$. In particular, $\Psi\in\mathbb{C}^{M\times K}$ is a truncated change-of-basis matrix whose columns are the vectorized Fourier-Zernike modes expressed in the pixel basis (position representation), $\Omega\in\mathbb{C}^{K\times K}$ is the cross-talk matrix of the mode sorter, and $C\in\mathbb{R}^{K\times K}$ is a diagonal nulling matrix describing the rejection of the fundamental mode.
By the additive property of Poisson random variables, the complete statistical model for the synthetic measurement is,
@@
where $\Lambda_{0}=N\lambda_{0}$ and $\Lambda_{B}=N\lambda_{B}$ are the effective photon fluxes for the star-planet system and the background respectively, with $N=N_{s}+1$ denoting the total number of measurement realizations. The photon distribution at the detector plane induced by the star-planet system is,